### v1

@@ 4.2 Finite-difference and finite-volume schemes
To capture the magnetar scenario in detail, numerical codes need to tackle a substantially more challenging regime. In Viganò et al (2012), a novel approach making use of the well-know High-Resolution Shock-Capturing (HRSC) techniques (Toro, 1997), designed to handle shocks in hydrodynamics and MHD, was proposed. These techniques have been successfully applied to a range of problems, from a simple 1D Burgers equation to complex ideal MHD problems (Antón et al, 2006; Giacomazzo and Rezzolla, 2007; Cerdá-Durán et al, 2008), avoiding the appearance of spurious oscillations near discontinuities. We refer to Mart´ı and Müller (2015) for a general review on grid-based methods and to Balsara (2017) for a review on finite-volume methods, applied to other astrophysical scenarios. Let us review some of the main characteristics of these methods, of particular interest in our problem.
@@ 4.2.1 Conservation form and staggered grids
With this definition, the divergence-preserving character of the methods using the conservation form to advance $\overline{B}_{\alpha}$ in time becomes evident: taking the time derivative of Eq. (42), and using Eq. (41), every edge contributes twice (one per each face) with opposite signs, so that all discrete contributions to the divergence time evolution exactly cancel out. Thus, by construction, the divergence condition is preserved to machine error for any divergence-free initial data. Examples of applications of such methods can be found, among many others, in Tóth (2000); Viganò et al (2012); Balsara and Dumbser (2015).
@@ 4.2.2 Divergence cleaning methods in finite-difference schemes
An algorithm built on a staggered grid can be designed to preserve the divergence constraint by construction, but the different allocation of variables makes its implementation relatively complex, particularly in 3D problems and with the inclusion of quadratic and cubic terms in the electric field. Among alternative formulations that have recently gained popularity, and can also handle many MHD-like problems, a relatively simple option is the family of divergence-cleaning schemes built on standard grids (all components of the fields are defined and evolved at every grid node). A popular divergence-cleaning method (Dedner et al, 2002), extensively used in MHD, consists in the extension of the system of equations as follows:
$\frac{1}{c}\frac{\partial\vec{B}}{\partial t}+\vec{\nabla}\times\vec{E}+\vec{\nabla}\chi=0~,$
$\frac{\partial\chi}{\partial t}+c_{h}^{2}\vec{\nabla}\cdot\vec{B}=-\gamma\chi~,$ (43)
@@ 4.2.4 Cell reconstruction and high-order accuracy
Other popular higher order reconstruction, are PPM (Colella and Woodward, 1984), PHM (Donat and Marquina, 1996), MP5 (Suresh and Huynh, 1997), the FDOC families (Bona et al, 2009), or the Weighted-Essentially-Non-Oscillatory (WENO) reconstructions (Jiang and Shu, 1996; Shu, 1998; Yamaleev and Carpenter, 2009; Balsara, 2017). In Viganò et al (2019) they presented and thoroughly tested a two-step method consisting of the reconstruction with WENO methods of a combination of fluxes and fields at each node, known as flux-splitting (Shu, 1998). This reconstruction scheme does not require the characteristic decomposition of the system of equations (i.e., the full spectrum of characteristic velocities) and, at the lowest order of reconstruction, their flux formula reduces to the popular and robust Local-Lax–Friedrichs flux (Toro, 1997).
@@ 4.3 Axis Singularities in 3D and Cubed-Sphere Grid
Figure 11: Figure adopted from Dehman et al (2023c). Exploded, cubed view of the patches (Ronchi et al, 1996). Each patch is identical and is described by the coordinates $\xi$ and $\eta$, both spanning the range $\pi/4;\pi/4]$. In the exploded view $\xi$ and $\eta$ grow to the right and upward, respectively, for all patches (only patch I is explicitly drawn here). Arrows identify the 12 edges between patches. The coordinate values $(\xi,\eta)$ of the corners for each of these patches are written in the bottom part as well.
@@
Common strategies to mitigate this problem include employing pseudo-spectral methods in the angular directions (e.g., MaGIc; Wicht 2002, Parody; Gourgouliatos et al 2016), excising the axis from the computational domain (e.g., the Pencil Code [^7] https://github.com/pencil-code in spherical geometry; Pencil Code Collaboration et al 2021), or adopting alternative coordinate systems such as Yin–Yang grids (Kageyama and Sato, 2004). In the newly developed MATINS code (Dehman et al, 2023c, b; Ascenzi et al, 2024; Dehman and Pons, 2025), the cubed-sphere coordinate system was implemented to overcome the axis singularity problem. Originally introduced by Ronchi et al (1996), this formalism keeps the radial direction as one coordinate, similar to spherical coordinates, while decomposing the volume into a stack of concentric radial layers. As illustrated in Fig. 11, each layer is covered by six non-overlapping patches. These patches can be visualized as the inflated faces of a cube expanded to a spherical surface. Unlike the Yin–Yang grid (Kageyama and Sato, 2004), the cubed-sphere has no overlap: its six patches meet edge-to-edge along great-circle arcs (Ronchi et al, 1996).
In MATINS, the cubed-sphere coordinates are prescribed as in the original work by Ronchi et al (1996), with the only difference given by the metric factors in the radial direction, according to Schwarzschild interior metric, obtained from a TOV solution. In this framework, each patch has the same functional form for the metric in terms of two coordinates mapping the spherical surface portion ($\xi,\eta$). The radial unit vector $\hat{e}_{r}$ is orthogonal to each spherical surface, but $\hat{e}_{\xi}$ and $\hat{e}_{\eta}$ are generally not orthogonal to each other, with the degree of skew varying across the patch. This non-orthogonality is evident in the spatial part of the metric tensor with coordinates ($r,\xi,\eta$): it contains non-zero off-diagonal terms in the ($\xi,\eta$) sub-block:
@@
Here, $X$, $Y,$ $C,$ and $D$ are auxiliary variables of the cubed-sphere grid. Further details or the explicit transformations between cubed-sphere, spherical, and Cartesian coordinates can be found in Ronchi et al (1996) or Appendix  A of Dehman et al (2023c).
@@
Figure 12: Figure adopted from Dehman et al (2023c). Schematic view of two contiguous equatorial blocks, e.g., patch I (black) and patch II (blue), and the ghosts cells of patch I (endpoints of the red dashes). The view is centered on the common vertical boundary line. The pseudo-horizontal coordinates $\xi$ of the ghost points of one grid, i.e., patch I, coincide with the second points along the $\xi$ coordinates of the last one interior grid points of the contiguous block, i.e., patch II. The ghost points are traced by the red line, and the values of the fields along the pseudo-vertical coordinate, $\eta$, are obtained by interpolations among the adjacent patch points (blue letters). Note that for other pairs of patches, the correspondence of coordinates may be less trivial (see Table 1 of Dehman et al (2023c)). A sketch of a centered discretized circulation which extends twice the size of the cell (once per each side around a central point $(i,j,k)$) is displayed on the left hand side of this plot, in red. The circulation shown here is applied to calculate the radial component of the curl operator for a given vector A, such as $(\vec{\nabla}\times\vec{A})^{r}$.
When evaluating derivatives near patch edges or corners, field values are needed at locations that lie in the coordinate system of adjacent patches (see Fig.12). In MATINS, this is handled by adding a single layer of ghost cells around each patch, which is sufficient for a second-order, centered finite-volume scheme. The field components in these ghost cells are obtained by interpolating neighbouring-patch data expressed in the local coordinates of the patch. Once these components are retrieved, Jacobian transformations are applied to convert them from the neighbour coordinate system to that of the patch where the derivative is actually being evaluated. As shown in Fig. 12, a ghost grid line in one patch coincides with an interior grid line in the neighbouring patch, requiring only a one-dimensional interpolation along the relevant angular coordinate. Because $\xi$ and $\eta$ share the same spacing ($d\xi=d\eta=\Delta$) and range $\pi/4,\pi/4]$, this procedure applies in both vertical and horizontal directions. Further details on patch-interface handling can be found in Dehman et al (2023c) and Appendix A of Dehman (2024).
@@ 4.4 Courant condition and time advance
$\Delta t_{h}\approx\frac{4\pi en_{c}L~\Delta l}{cB}~,$ (48)
@@
$\Delta t_{a}\approx\frac{4\pi L~\Delta l}{cf_{a}B^{2}}~,$ (50)
@@
$\Delta t=k_{c}\,\mbox{min}\left[\Delta t_{h},\Delta t_{5},\Delta t_{a}\right]~,$ (51)
@@
Other stabilizing techniques introduced in O’Sullivan and Downes (2006) for the time advance of the non-linear terms are used in González-Morales et al (2018). These methods, namely the Super Time-Stepping and the Hall Diffusion Schemes, allow the code to maintain stability and efficiently speed up the time evolution when the ambipolar or the Hall term dominates. Another common technique is the use of high-order dissipation (also called hyper-resistivity; Huba 2003), or a predictor-corrector step advancing alternatively different field components.
Viganò et al (2012) used a particularly simple method that significantly improves the stability of the scheme in spherical coordinates. Their procedure to advance the solution from $t_{n}$ to $t_{n+1}=t_{n}+\Delta t$ can be summarized as follows:
@@
In Tóth et al (2008), the authors discussed that such a two-stage formulation is equivalent to introducing a fourth-order hyper-resistivity. Since the toroidal component is advanced first, it follows that the hyper-resistive correction only acts on the evolution of the poloidal components. In Viganò et al (2012) it was also shown that the additional correction given by $\vec{E}_{t}^{\star}$ contains higher-order spatial derivatives and scales with $(\Delta t)^{2}$, which is characteristic of hyper-resistive terms. They found a significant improvement in the stability of the method when comparing a fully explicit algorithm with the two-steps method, allowing to work with $k_{c}\approx 10^{-2}-10^{-1}$.
In the finite-difference schemes of Viganò et al (2019), the authors used a fourth-order Runge–Kutta scheme and found that the instabilities are especially significant when using fifth-order-accurate methods for the flux reconstruction (i.e. WENO5), which needed to be combined with the application of artificial Kreiss–Oliger dissipation along each coordinate direction (Calabrese et al, 2004). A sixth-order derivative dissipation operator has a similar stabilizing effect, filtering the high-frequency modes which cannot be accurately resolved by the numerical grid, at the cost of a potential loss of accuracy (Viganò et al, 2019). For this reason, they recommend using third-order schemes that do not require any additional artificial Kreiss–Oliger dissipation. The typical Courant factors used were again quite low, $k_{c}\approx 10^{-2}-10^{-1}$.
@@ 5 Magnetosphere-interior coupling and rotational evolution
The dynamics of the magnetosphere is essentially governed by the electro-magnetic field, since the plasma pressure and inertia are negligible. Therefore, a suitable approximation is that the large-scale magnetospheric structure follows force-free configurations, where electric and magnetic forces on the plasma are perfectly balanced, as in a perfect conductor (see Cerutti and Beloborodov (2017); Philippov and Kramer (2022) for comprehensive reviews on electrodynamics of pulsar magnetospheres).
@@
For magnetar conditions, rotational effects in the magnetospheric region close to the star can be safely ignored and the force-free condition reduces to $\vec{j}\times\vec{B}=0$: the electric currents flow parallel to the magnetic field lines that they sustain [^8] Since $\vec{j}\propto\vec{\nabla}\times\vec{B}$, a force-free magnetic field is a also Beltrami vector field. Among possible solutions, the simplest and most commonly used is the current-free or potential solution, $\vec{j}=0$, which also applies in vacuum. Matching the interior magnetic field to a magnetospheric potential field effectively prevents current from escaping or entering the star. However, a non-zero Poynting flux across the boundary enables magnetic energy exchange between the two regions, though magnetic helicity remains conserved.
@@
Although this simple current-free solution (adopted by the majority of works) serves as a reasonable first approximation, developing more realistic models requires more general solutions. In particular, stable electrical currents can flow within regions of closed magnetic field lines, much like the coronal loops observed on the Sun. These current systems can persist for relatively long timescales, from months to decades (Beloborodov, 2009), likely sustained by the interior dynamics. There is indirect observational evidence of such currents in some magnetars, where the presence of a plasma much denser than the Goldreich-Julian value has been inferred. Soft X-ray photons emitted from the star surface are up-scattered to higher energy (Lyutikov and Gavriil, 2006; Rea et al, 2008; Beloborodov, 2013) through resonant Compton processes, resulting in the observed spectra.
@@
Equilibrium solutions for force-free twisted magnetospheres in the context of magnetars have been investigated in several studies (Fujisawa and Kisaka, 2014; Glampedakis et al, 2014; Pili et al, 2015; Akgün et al, 2016; Kojima, 2017). However, the interior evolution can sometimes lead to configurations that cannot be smoothly matched to a force-free exterior. This mismatch results in discontinuities in the tangential magnetic field components at the surface, corresponding to current sheets, which may introduce numerical instabilities.
@@
Moreover, solving a fully consistent 2D or 3D problem at each timestep in global simulations is computationally expensive, primarily due to the demands of elliptic solvers required to solve exactly the exterior force-free condition. To address these challenges, a novel approach has recently been proposed (Urbán et al, 2023; Stefanou et al, 2023b), exploring the use of PINNs for modeling the magnetic field evolution inside a NS coupled to a force-free magnetosphere. This method offers a promising alternative to traditional techniques. Initial results show that PINN-based solutions are accurate, robust, and numerically stable. Notably, Stefanou et al (2023b) found the computational cost to be over an order of magnitude lower than that of comparable simulations using conventional methods, and there is plenty of room for improvement (Urbán et al, 2025). These findings opened the door to extensions to fully 3D problems (Stefanou et al, 2025), where implementing generalized boundary conditions becomes even more computationally demanding.

### v2

@@ 4.2 Finite-difference and finite-volume schemes
To capture the magnetar scenario in detail, numerical codes need to tackle a substantially more challenging regime. In Viganò et al. (2012), a novel approach making use of the well-know High-Resolution Shock-Capturing (HRSC) techniques (Toro, 1997), designed to handle shocks in hydrodynamics and MHD, was proposed. These techniques have been successfully applied to a range of problems, from a simple 1D Burgers equation to complex ideal MHD problems (Antón et al., 2006; Giacomazzo and Rezzolla, 2007; Cerdá-Durán et al., 2008), avoiding the appearance of spurious oscillations near discontinuities. We refer to Martí and Müller (2015) for a general review on grid-based methods and to Balsara (2017) for a review on finite-volume methods, applied to other astrophysical scenarios. Let us review some of the main characteristics of these methods, of particular interest in our problem.
@@ 4.2.1 Conservation form and staggered grids
With this definition, the divergence-preserving character of the methods using the conservation form to advance $\overline{B}_{\alpha}$ in time becomes evident: taking the time derivative of Eq. (42), and using Eq. (41), every edge contributes twice (one per each face) with opposite signs, so that all discrete contributions to the divergence time evolution exactly cancel out. Thus, by construction, the divergence condition is preserved to machine error for any divergence-free initial data. Examples of applications of such methods can be found, among many others, in Tóth (2000); Viganò et al. (2012); Balsara and Dumbser (2015).
@@ 4.2.2 Divergence cleaning methods in finite-difference schemes
An algorithm built on a staggered grid can be designed to preserve the divergence constraint by construction, but the different allocation of variables makes its implementation relatively complex, particularly in 3D problems and with the inclusion of quadratic and cubic terms in the electric field. Among alternative formulations that have recently gained popularity, and can also handle many MHD-like problems, a relatively simple option is the family of divergence-cleaning schemes built on standard grids (all components of the fields are defined and evolved at every grid node). A popular divergence-cleaning method (Dedner et al., 2002), extensively used in MHD, consists in the extension of the system of equations as follows:
$\frac{1}{c}\frac{\partial\vec{B}}{\partial t}+\vec{\nabla}\times\vec{E}+\vec{\nabla}\chi=0\,$,
$\frac{\partial\chi}{\partial t}+c_{h}^{2}\vec{\nabla}\cdot\vec{B}=-\gamma\chi\,$, (43)
@@ 4.2.4 Cell reconstruction and high-order accuracy
Other popular higher order reconstruction, are PPM (Colella and Woodward, 1984), PHM (Donat and Marquina, 1996), MP5 (Suresh and Huynh, 1997), the FDOC families (Bona et al., 2009), or the Weighted-Essentially-Non-Oscillatory (WENO) reconstructions (Jiang and Shu, 1996; Shu, 1998; Yamaleev and Carpenter, 2009; Balsara, 2017). In Viganò et al. (2019) they presented and thoroughly tested a two-step method consisting of the reconstruction with WENO methods of a combination of fluxes and fields at each node, known as flux-splitting (Shu, 1998). This reconstruction scheme does not require the characteristic decomposition of the system of equations (i.e., the full spectrum of characteristic velocities) and, at the lowest order of reconstruction, their flux formula reduces to the popular and robust Local-Lax–Friedrichs flux (Toro, 1997).
@@ 4.3 Axis Singularities in 3D and Cubed-Sphere Grid
Figure 11: Exploded, cubed view of the patches (Ronchi et al., 1996). Each patch is identical and is described by the coordinates $\xi$ and $\eta$, both spanning the range $\pi/4;\pi/4]$. In the exploded view $\xi$ and $\eta$ grow to the right and upward, respectively, for all patches (only patch I is explicitly drawn here). Arrows identify the 12 edges between patches. The coordinate values $(\xi,\eta)$ of the corners for each of these patches are written in the bottom part as well. Image reproduced with permission from Dehman et al. (2023c), copyright by the author(s).
@@
Common strategies to mitigate this problem include employing pseudo-spectral methods in the angular directions (e.g., MaGIc, Wicht 2002; Parody, Gourgouliatos et al. 2016), excising the axis from the computational domain (e.g., the Pencil Code [^7] https://github.com/pencil-code in spherical geometry, Pencil Code Collaboration et al. 2021), or adopting alternative coordinate systems such as Yin–Yang grids (Kageyama and Sato, 2004). In the newly developed MATINS  [^8] https://github.com/ice-csic-astroexotic/MATINS code (Dehman et al., 2023c, b; Ascenzi et al., 2024; Dehman and Pons, 2025), the cubed-sphere coordinate system was implemented to overcome the axis singularity problem. Originally introduced by Ronchi et al. (1996), this formalism keeps the radial direction as one coordinate, similar to spherical coordinates, while decomposing the volume into a stack of concentric radial layers. As illustrated in Fig. 11, each layer is covered by six non-overlapping patches. These patches can be visualized as the inflated faces of a cube expanded to a spherical surface. Unlike the Yin–Yang grid (Kageyama and Sato, 2004), the cubed-sphere has no overlap: its six patches meet edge-to-edge along great-circle arcs (Ronchi et al., 1996).
In MATINS, the cubed-sphere coordinates are prescribed as in the original work by Ronchi et al. (1996), with the only difference given by the metric factors in the radial direction, according to Schwarzschild interior metric, obtained from a TOV solution. In this framework, each patch has the same functional form for the metric in terms of two coordinates mapping the spherical surface portion ($\xi,\eta$). The radial unit vector $\hat{e}_{r}$ is orthogonal to each spherical surface, but $\hat{e}_{\xi}$ and $\hat{e}_{\eta}$ are generally not orthogonal to each other, with the degree of skew varying across the patch. This non-orthogonality is evident in the spatial part of the metric tensor with coordinates ($r,\xi,\eta$): it contains non-zero off-diagonal terms in the ($\xi,\eta$) sub-block:
@@
Here, $X$, $Y,$ $C,$ and $D$ are auxiliary variables of the cubed-sphere grid. Further details or the explicit transformations between cubed-sphere, spherical, and Cartesian coordinates can be found in Ronchi et al. (1996) or Appendix A of Dehman et al. (2023c).
@@
Figure 12: Schematic view of two contiguous equatorial blocks, e.g., patch I (black) and patch II (blue), and the ghosts cells of patch I (endpoints of the red dashes). The view is centered on the common vertical boundary line. The pseudo-horizontal coordinates $\xi$ of the ghost points of one grid, i.e., patch I, coincide with the second points along the $\xi$ coordinates of the last one interior grid points of the contiguous block, i.e., patch II. The ghost points are traced by the red line, and the values of the fields along the pseudo-vertical coordinate, $\eta$, are obtained by interpolations among the adjacent patch points (blue letters). Note that for other pairs of patches, the correspondence of coordinates may be less trivial (see Table 1 of Dehman et al. 2023c). A sketch of a centered discretized circulation which extends twice the size of the cell (once per each side around a central point $(i,j,k)$) is displayed on the left hand side of this plot, in red. The circulation shown here is applied to calculate the radial component of the curl operator for a given vector A, such as $(\vec{\nabla}\times\vec{A})^{r}$. Image reproduced with permission from Dehman et al. (2023c), copyright by the author(s).
When evaluating derivatives near patch edges or corners, field values are needed at locations that lie in the coordinate system of adjacent patches (see Fig.12). In MATINS, this is handled by adding a single layer of ghost cells around each patch, which is sufficient for a second-order, centered finite-volume scheme. The field components in these ghost cells are obtained by interpolating neighbouring-patch data expressed in the local coordinates of the patch. Once these components are retrieved, Jacobian transformations are applied to convert them from the neighbour coordinate system to that of the patch where the derivative is actually being evaluated. As shown in Fig. 12, a ghost grid line in one patch coincides with an interior grid line in the neighbouring patch, requiring only a one-dimensional interpolation along the relevant angular coordinate. Because $\xi$ and $\eta$ share the same spacing ($d\xi=d\eta=\Delta$) and range $\pi/4,\pi/4]$, this procedure applies in both vertical and horizontal directions. Further details on patch-interface handling can be found in Dehman et al. (2023c) and Appendix A of Dehman (2024).
@@ 4.4 Courant condition and time advance
$\Delta t_{h}\approx\frac{4\pi en_{c}L~\Delta l}{cB}\,$, (48)
@@
$\Delta t_{a}\approx\frac{4\pi L~\Delta l}{cf_{a}B^{2}}\,$, (50)
@@
$\Delta t=k_{c}\,\mbox{min}\left[\Delta t_{h},\Delta t_{5},\Delta t_{a}\right]\,$, (51)
@@
Other stabilizing techniques introduced in O’Sullivan and Downes (2006) for the time advance of the non-linear terms are used in González-Morales et al. (2018). These methods, namely the Super Time-Stepping and the Hall Diffusion Schemes, allow the code to maintain stability and efficiently speed up the time evolution when the ambipolar or the Hall term dominates. Another common technique is the use of high-order dissipation (also called hyper-resistivity; Huba 2003), or a predictor-corrector step advancing alternatively different field components.
Viganò et al. (2012) used a particularly simple method that significantly improves the stability of the scheme in spherical coordinates. Their procedure to advance the solution from $t_{n}$ to $t_{n+1}=t_{n}+\Delta t$ can be summarized as follows:
@@
In Tóth et al. (2008), the authors discussed that such a two-stage formulation is equivalent to introducing a fourth-order hyper-resistivity. Since the toroidal component is advanced first, it follows that the hyper-resistive correction only acts on the evolution of the poloidal components. In Viganò et al. (2012) it was also shown that the additional correction given by $\vec{E}_{t}^{\star}$ contains higher-order spatial derivatives and scales with $(\Delta t)^{2}$, which is characteristic of hyper-resistive terms. They found a significant improvement in the stability of the method when comparing a fully explicit algorithm with the two-steps method, allowing to work with $k_{c}\approx 10^{-2}-10^{-1}$.
In the finite-difference schemes of Viganò et al. (2019), the authors used a fourth-order Runge–Kutta scheme and found that the instabilities are especially significant when using fifth-order-accurate methods for the flux reconstruction (i.e. WENO5), which needed to be combined with the application of artificial Kreiss–Oliger dissipation along each coordinate direction (Calabrese et al., 2004). A sixth-order derivative dissipation operator has a similar stabilizing effect, filtering the high-frequency modes which cannot be accurately resolved by the numerical grid, at the cost of a potential loss of accuracy (Viganò et al., 2019). For this reason, they recommend using third-order schemes that do not require any additional artificial Kreiss–Oliger dissipation. The typical Courant factors used were again quite low, $k_{c}\approx 10^{-2}-10^{-1}$.
@@ 5 Magnetosphere-interior coupling and rotational evolution
The dynamics of the magnetosphere is essentially governed by the electro-magnetic field, since the plasma pressure and inertia are negligible. Therefore, a suitable approximation is that the large-scale magnetospheric structure follows force-free configurations, where electric and magnetic forces on the plasma are perfectly balanced, as in a perfect conductor (see Cerutti and Beloborodov 2017; Philippov and Kramer 2022 for comprehensive reviews on electrodynamics of pulsar magnetospheres).
@@
For magnetar conditions, rotational effects in the magnetospheric region close to the star can be safely ignored and the force-free condition reduces to $\vec{j}\times\vec{B}=0$: the electric currents flow parallel to the magnetic field lines that they sustain. [^9] Since $\vec{j}\propto\vec{\nabla}\times\vec{B}$, a force-free magnetic field is a also Beltrami vector field. Among possible solutions, the simplest and most commonly used is the current-free or potential solution, $\vec{j}=0$, which also applies in vacuum. Matching the interior magnetic field to a magnetospheric potential field effectively prevents current from escaping or entering the star. However, a non-zero Poynting flux across the boundary enables magnetic energy exchange between the two regions, though magnetic helicity remains conserved.
@@
Although this simple current-free solution (adopted by the majority of works) serves as a reasonable first approximation, developing more realistic models requires more general solutions. In particular, stable electrical currents can flow within regions of closed magnetic field lines, much like the coronal loops observed on the Sun. These current systems can persist for relatively long timescales, from months to decades (Beloborodov, 2009), likely sustained by the interior dynamics. There is indirect observational evidence of such currents in some magnetars, where the presence of a plasma much denser than the Goldreich-Julian value has been inferred. Soft X-ray photons emitted from the star surface are up-scattered to higher energy (Lyutikov and Gavriil, 2006; Rea et al., 2008; Beloborodov, 2013) through resonant Compton processes, resulting in the observed spectra.
@@
Equilibrium solutions for force-free twisted magnetospheres in the context of magnetars have been investigated in several studies (Fujisawa and Kisaka, 2014; Glampedakis et al., 2014; Pili et al., 2015; Akgün et al., 2016; Kojima, 2017). However, the interior evolution can sometimes lead to configurations that cannot be smoothly matched to a force-free exterior. This mismatch results in discontinuities in the tangential magnetic field components at the surface, corresponding to current sheets, which may introduce numerical instabilities.
@@
Moreover, solving a fully consistent 2D or 3D problem at each timestep in global simulations is computationally expensive, primarily due to the demands of elliptic solvers required to solve exactly the exterior force-free condition. To address these challenges, a novel approach has recently been proposed (Urbán et al., 2023; Stefanou et al., 2023b), exploring the use of PINNs for modeling the magnetic field evolution inside a NS coupled to a force-free magnetosphere. This method offers a promising alternative to traditional techniques. Initial results show that PINN-based solutions are accurate, robust, and numerically stable. Notably, Stefanou et al. (2023b) found the computational cost to be over an order of magnitude lower than that of comparable simulations using conventional methods, and there is plenty of room for improvement (Urbán et al., 2025). These findings opened the door to extensions to fully 3D problems (Stefanou et al., 2025), where implementing generalized boundary conditions becomes even more computationally demanding.